OpenStudy (anonymous):

Verify the identity. Cot[theta - pi/2] = -tan theta

OpenStudy (anonymous):

\[Cot \left( \theta - \frac{ \pi }{ 2 } \right) = -\tan \theta\]

OpenStudy (anonymous):

@oleg3321

OpenStudy (anonymous):

I have to verify the identity and show my work

imqwerty (imqwerty):

we know that cot(-theta) = -cot(theta) we also know that cot(pi/2-theta) = tan(theta) multiply both sides by -1 we get -- -cot(pi/2-theta)=-tan(theta) we can write -cot(pi/2-theta) as cot(theta-pi/2) so nw we hav cot(theta-pi/2)= -tan(theta)

OpenStudy (anonymous):

ok so where do i go from there? Im literally lost and I need to show my work

OpenStudy (anonymous):

Or is that the answer? @imqwerty

OpenStudy (anonymous):

Anyone?

imqwerty (imqwerty):

wait :)

OpenStudy (anonymous):

ok lol sorry

imqwerty (imqwerty):

don't get lost :) that is the answer @EllenJaz17

OpenStudy (anonymous):

oh ok, lol thank you!

imqwerty (imqwerty):

ur welcome :) @EllenJaz17